# Dongyue Lu1, Lingdong Kong1, Gim Hee Lee1, Camille Simon Chane2, Wei Tsang Ooi1

### v1

@@ A.2 Baselines
  • Event-Only Methods. We include state-of-the-art event-only object detectors, namely RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), LEOD (Wu et al., 2024), and SSM (Zubić et al., 2024), which are originally trained on event-only datasets like Gen1 (De Tournemire et al., 2020) and 1Mpx (Perot et al., 2020). To ensure a fair comparison, we retrain these methods on the DSEC-Det dataset following their respective training protocols.
  • Event-Frame Fusion Methods. For event-frame fusion methods on DSEC-Det, we include DAGr, as it has been evaluated on this dataset. We report the scores of DAGr (Gehrig & Scaramuzza, 2024) from the original paper to ensure consistency and fairness. For the DSEC-Detection and DSEC-MOD datasets, we train our model following the standard training and evaluation settings, and compare it against state-of-the-art methods CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023), as reported in their respective papers. For other methods evaluated on DSEC-Detection and DSEC-MOD, we reference the results reported in the CAFR and RENet papers, respectively. Since DAGr’s training code is not publicly available, we could not reproduce its results on DSEC-Detection and DSEC-MOD.
These comparisons ensure a fair and comprehensive evaluation while adhering to resource and code availability constraints.
@@ A.3 Evaluation Metrics
In addition to mAP, we also report the following metrics from the COCO evaluation protocol (Lin et al., 2014):
@@
A.4 Training & Inference Details
We train our models using mixed precision to optimize both memory efficiency and training speed. The training process spans 100,000 iterations, utilizing the ADAM optimizer (Kingma, 2014) with a OneCycle learning rate schedule (Smith & Topin, 2019), which gradually decays from a peak learning rate to enhance convergence.
Consistent with RVT (Gehrig & Scaramuzza, 2023), we employ a mixed batching strategy to balance computational efficiency and memory usage. Specifically:
Our training process utilize the YOLOX framework (Zheng et al., 2021), a versatile object detection framework known for its efficient and high-performing architecture. We employ a multi-component loss function to optimize our model effectively:
To ensure stable training, these loss components are averaged across both the batch and sequence length at each optimization step. This averaging process helps to reduce variance during training and facilitates smoother convergence of the model parameters.
Training Configuration. The training is conducted with a batch size of 8, which provides an optimal balance between efficient GPU utilization and memory requirements. Each training sample contains a sequence length of 11 frames, allowing the model to learn temporal dependencies effectively. The frame backbone’s weights are initialized using pre-trained ResNet and the event backbone’s weights are initialized using pre-trained RVT. The learning rate is set to $1\times 10^{-4}$, following a OneCycle learning rate schedule that allows efficient exploration of the learning space and helps in achieving faster convergence.
Hardware & Training Time. All training experiments are carried out on two NVIDIA RTX A5000 GPUs, each with 24GB of memory, providing the computational resources necessary for handling the high-resolution event data and RGB frames. The complete training process, including all iterations and model optimization, takes approximately one day, demonstrating the efficiency of our implementation in terms of both training speed and resource utilization.
Table 7: Comparative efficiency analysis of state-of-the-art event camera detectors on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024), comparing both event-only and event-frame fusion methods. This table reports inference times at various frequencies, measured in milliseconds (ms).
Modality Method Param Frequency (Hz)
(M) $\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30}$ $\mathbf{36}$ $\mathbf{45}$ $\mathbf{60}$ $\mathbf{90}$ $\mathbf{180}$ $\mathbf{200}$
E RVT $18.5$ $9.20$ $8.67$ $8.35$ $7.93$ $7.61$ $7.51$ $7.19$ $6.77$ $6.34$
SAST 18.9 $14.06$ $13.11$ $12.68$ $12.37$ $11.95$ $11.63$ $11.52$ $11.10$ $10.36$
SSM 18.2 $8.79$ $8.26$ $8.08$ $7.71$ $7.55$ $7.30$ $6.90$ $6.54$ $6.12$
E + F DAGr-50 $34.6$ $73.35$ $65.73$ $60.02$ $55.11$ $51.00$ $48.00$ $45.29$ $43.89$ $37.58$
FlexEvent $45.4$ $14.27$ $13.53$ $13.32$ $13.00$ $12.79$ $12.58$ $12.47$ $12.37$ $12.12$
Table 8: The complete results of the ablation study (cf. Tab. 4) of different components in the FlexEvent framework. FlexFuser denotes the adaptive event-frame fusion module (cf. Sec.3.2). FAL denotes the frequency-adaptive learning module (cf. Sec. 3.3). The reported results are the mAP, AP50, AP75, APS, APM, and APL scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FAL. The best and 2nd best scores of each metric are highlighted in bold and underline, respectively.
Modality FAL FlexFuser Frequency (Hz) mAP AP50 AP75 APS APM APL
@@ A.4 Training & Inference Details
$\blacklozenge$ $45.0$ $57.2\%$ $75.5\%$ $65.4\%$ $51.6\%$ $\mathbf{68.2\%}$ $\mathbf{87.5\%}$
@@
Table 9: The complete results of the ablation study (cf. Tab. 5) of different hyperparameter configurations in the FlexEvent framework. $\mathbf{\tau}^{\text{car}}$, $\mathbf{\tau}^{\text{ped}}$ denotes the confidence threshold for car and pedestrian, respectively. $\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length. The reported results are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). The best and 2nd best scores of each metric from each hyperparameter configuration are highlighted in bold and underline, respectively.
@@
$0.6$ $0.3$ $8$ $0.6$ $20.0$ $56.3\%$ $77.2\%$ $64.9\%$ $50.4\%$ $64.1\%$ $83.3\%$
$0.6$ $0.3$ $8$ $0.6$ $27.5$ $58.5\%$ $78.3\%$ $68.1\%$ $52.5\%$ $66.8\%$ $84.8\%$
$0.6$ $0.3$ $8$ $0.6$ $30.0$ $58.8\%$ $78.5\%$ $68.7\%$ $52.6\%$ $67.9\%$ $85.7\%$
$0.6$ $0.3$ $8$ $0.6$ $36.0$ $\mathbf{59.1\%}$ $\mathbf{78.8\%}$ $\mathbf{69.0\%}$ $\mathbf{52.7\%}$ $68.8\%$ $\mathbf{86.5\%}$
$0.6$ $0.3$ $8$ $0.6$ $45.0$ $58.8\%$ $78.2\%$ $68.6\%$ $52.3\%$ $68.8\%$ $86.1\%$
$0.6$ $0.3$ $8$ $0.6$ $60.0$ $58.4\%$ $77.9\%$ $67.5\%$ $51.3\%$ $\mathbf{69.6\%}$ $85.5\%$
$0.6$ $0.3$ $8$ $0.6$ $90.0$ $56.2\%$ $76.6\%$ $64.8\%$ $48.5\%$ $68.3\%$ $84.7\%$
$0.6$ $0.3$ $8$ $0.6$ $180.0$ $51.2\%$ $71.9\%$ $56.3\%$ $42.6\%$ $64.2\%$ $82.9\%$
$0.6$ $0.3$ $6$ $0.6$ $20.0$ $57.3\%$ $80.0\%$ $65.2\%$ $51.2\%$ $65.8\%$ $84.1\%$
$0.6$ $0.3$ $6$ $0.6$ $27.5$ $59.4\%$ $81.3\%$ $68.5\%$ $53.4\%$ $68.8\%$ $\mathbf{85.7\%}$
$0.6$ $0.3$ $6$ $0.6$ $30.0$ $59.7\%$ $\mathbf{81.7\%}$ $\mathbf{69.0\%}$ $\mathbf{53.7\%}$ $69.0\%$ $85.3\%$
$0.6$ $0.3$ $6$ $0.6$ $36.0$ $\mathbf{59.9\%}$ $81.4\%$ $\mathbf{69.0\%}$ $53.6\%$ $\mathbf{69.8\%}$ $\mathbf{85.7\%}$
$0.6$ $0.3$ $6$ $0.6$ $45.0$ $59.3\%$ $80.5\%$ $67.9\%$ $52.8\%$ $69.4\%$ $85.6\%$
$0.6$ $0.3$ $6$ $0.6$ $60.0$ $58.5\%$ $79.6\%$ $67.0\%$ $51.5\%$ $69.6\%$ $84.9\%$
$0.6$ $0.3$ $6$ $0.6$ $90.0$ $55.7\%$ $77.2\%$ $62.8\%$ $48.1\%$ $67.9\%$ $84.3\%$
$0.6$ $0.3$ $6$ $0.6$ $180.0$ $48.8\%$ $70.6\%$ $50.9\%$ $40.8\%$ $61.1\%$ $83.4\%$
$0.6$ $0.6$ $6$ $0.6$ $20.0$ $57.4\%$ $78.2\%$ $66.6\%$ $51.7\%$ $64.9\%$ $83.7\%$
$0.6$ $0.6$ $6$ $0.6$ $27.5$ $60.0\%$ $79.4\%$ $70.1\%$ $53.5\%$ $68.4\%$ $\mathbf{86.1\%}$
$0.6$ $0.6$ $6$ $0.6$ $30.0$ $60.0\%$ $\mathbf{79.7\%}$ $\mathbf{70.8\%}$ $\mathbf{53.6\%}$ $69.9\%$ $\mathbf{86.1\%}$
$0.6$ $0.6$ $6$ $0.6$ $36.0$ $\mathbf{60.1\%}$ $79.6\%$ $\mathbf{70.8\%}$ $53.2\%$ $70.3\%$ $85.7\%$
$0.6$ $0.6$ $6$ $0.6$ $45.0$ $59.5\%$ $79.0\%$ $69.5\%$ $52.5\%$ $70.8\%$ $85.3\%$
$0.6$ $0.6$ $6$ $0.6$ $60.0$ $58.8\%$ $78.5\%$ $69.0\%$ $51.1\%$ $\mathbf{71.1\%}$ $85.3\%$
$0.6$ $0.6$ $6$ $0.6$ $90.0$ $56.5\%$ $76.5\%$ $65.4\%$ $48.2\%$ $70.1\%$ $83.8\%$
$0.6$ $0.6$ $6$ $0.6$ $180.0$ $50.9\%$ $71.4\%$ $56.2\%$ $41.6\%$ $65.4\%$ $82.9\%$
$0.8$ $0.8$ $6$ $0.6$ $20.0$ $56.6\%$ $80.7\%$ $65.5\%$ $50.8\%$ $65.4\%$ $82.6\%$
$0.8$ $0.8$ $6$ $0.6$ $27.5$ $58.7\%$ $81.9\%$ $68.9\%$ $\mathbf{52.8\%}$ $68.6\%$ $84.9\%$
$0.8$ $0.8$ $6$ $0.6$ $30.0$ $\mathbf{59.1\%}$ $\mathbf{82.0\%}$ $\mathbf{69.2\%}$ $52.7\%$ $69.5\%$ $84.8\%$
$0.8$ $0.8$ $6$ $0.6$ $36.0$ $58.9\%$ $81.7\%$ $68.8\%$ $52.6\%$ $\mathbf{69.9\%}$ $85.0\%$
$0.8$ $0.8$ $6$ $0.6$ $45.0$ $58.4\%$ $81.4\%$ $67.7\%$ $51.6\%$ $69.8\%$ $\mathbf{85.1\%}$
$0.8$ $0.8$ $6$ $0.6$ $60.0$ $57.4\%$ $80.1\%$ $66.6\%$ $50.2\%$ $\mathbf{69.9\%}$ $84.3\%$
$0.8$ $0.8$ $6$ $0.6$ $90.0$ $55.7\%$ $78.6\%$ $63.7\%$ $47.8\%$ $68.9\%$ $84.2\%$
$0.8$ $0.8$ $6$ $0.6$ $180.0$ $50.2\%$ $74.0\%$ $55.2\%$ $41.5\%$ $63.7\%$ $83.1\%$
Appendix B B Additional Quantitative Results
B.1 Comparison of Efficiency
We present a comparative analysis of inference times and parameter counts for the evaluated methods in Tab. 7. All experiments are conducted on an NVIDIA RTX A5000 24GB GPU paired with an AMD EPYC 9354P 32-Core Processor operating at 3.8 GHz. The results demonstrate that, despite having a higher parameter count, FlexEvent achieves inference times comparable to the event-only method SAST and significantly outperforms the event-frame fusion method DAGr in terms of speed. Additionally, FlexEvent consistently surpasses DAGr in inference time across all tested frequencies. These results underscore the efficiency and rapid performance of FlexEvent, highlighting its suitability for real-time applications.
@@ B.2 Complete Results of Ablation Study
B.3 Complete Results of Hyperparameter Searching
@@
Appendix C C Additional Qualitative Results
C.1 Visual Comparisons of Event Camera Detectors
We include additional qualitative assessments in Fig. 7, Fig. 8, and Fig. 9.
C.2 Visual Comparisons under Different Frequencies
We include additional qualitative assessments in Fig. 10, Fig. 11, and Fig. 12.
Appendix D D Potential Societal Impact & Limitations

### v3

@@ A.2 Baselines
  • Event-Only Methods. We include state-of-the-art event-only object detectors, namely RVT [18], SAST [39], LEOD [55], and SSM [70], which are originally trained on event-only datasets like Gen11 [10] and 11Mpx [40]. To ensure a fair comparison, we retrain these methods on the DSEC-Det [16] dataset following their respective training protocols.
  • Event-Frame Fusion Methods. For event-frame fusion methods on DSEC-Det [16], we include DAGr [16] and HDI-Former [31], as they have been evaluated on this dataset. We report the scores of DAGr and HDI-Former from the original paper to ensure consistency and fairness. For the DSEC-Detection [50] and DSEC-MOD [66] datasets, we train our model following the standard training and evaluation settings, and compare it against state-of-the-art methods CAFR [3] and RENet [66], as reported in their respective papers. For other methods evaluated on DSEC-Detection [50] and DSEC-MOD [66], we reference the results reported in the CAFR and RENet papers, respectively.
These comparisons among state-of-the-art works ensure a fair and comprehensive evaluation while adhering to resource and code availability constraints.
@@ A.3 Evaluation Metrics
In addition to mAP, we also report the following metrics from the COCO evaluation protocol [35]:
@@
Table 8: The complete results of the efficiency analysis of state-of-the-art event detectors on the test set of DSEC-Det [16], comparing both event-only and event-frame fusion methods. This table reports inference times at various frequencies, measured in milliseconds (ms).
Modal Method Param Operational Frequency
(M) $\mathbf{20.0}$ Hz $\mathbf{36}$ Hz $\mathbf{45}$ Hz $\mathbf{60}$ Hz $\mathbf{90}$ Hz $\mathbf{180}$ Hz $\mathbf{200}$ Hz
E RVT [18] $18.5$ $9.20$ ms $7.93$ ms $7.61$ ms $7.51$ ms $7.19$ ms $6.77$ ms $6.34$ ms
SAST [39] $18.9$ $14.06$ ms $12.37$ ms $11.95$ ms $11.63$ ms $11.52$ ms $11.10$ ms $10.36$ ms
SSM [70] $18.2$ $8.79$ ms $7.71$ ms $7.55$ ms $7.30$ ms $6.90$ ms $6.54$ ms $6.12$ ms
E + F DAGr-50 [16] $34.6$ $73.35$ ms $55.11$ ms $51.00$ ms $48.00$ ms $45.29$ ms $43.89$ ms $37.58$ ms
FlexEvent $45.4$ $14.27$ ms $13.00$ ms $12.79$ ms $12.58$ ms $12.47$ ms $12.37$ ms $12.12$ ms
Table 9: The complete ablation results of different components in the FlexEvent framework. FlexFuse denotes the adaptive event-frame fusion module. FlexTune denotes the FlexTune learning mechanism. The reported results are the mAP, AP50, AP75, APS, APM, and APL scores on the test set of DSEC-Det [16]. The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FlexTune. The best and 2nd best scores of each metric are highlighted in bold and underline.
Modality FlexTune FlexFuse Frequency (Hz) mAP AP50 AP75 APS APM APL
@@
$\blacklozenge$ $45.0$ $2\% 75.57$$5\%$ $65.4\%$ $51.6\%$ $\mathbf{68.2\%}$ $\mathbf{87.5\%}$
@@
Table 10: The complete ablation results of different hyperparameter configurations in the FlexEvent framework. $\mathbf{\tau}^{\text{car}}$, $\mathbf{\tau}^{\text{ped}}$ denotes the confidence threshold for car and pedestrian, respectively. $\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length. The reported results are the mAP scores on the test set of DSEC-Det [16]. The best and 2nd best scores of each metric from each hyperparameter configuration are highlighted in bold and underline.
@@
$0.6$ $0.3$ 88 $0.6$ $20.0$ $56.3\%$ $77.2\%$ $64.9\%$ $50.4\%$ $64.1\%$ $83.3\%$
$0.6$ $0.3$ 88 $0.6$ $27.5$ $58.5\%$ $78.3\%$ $68.1\%$ $52.5\%$ $66.8\%$ $84.8\%$
$0.6$ $0.3$ 88 $0.6$ $30.0$ $58.8\%$ $78.5\%$ $68.7\%$ $52.6\%$ $67.9\%$ $85.7\%$
$0.6$ $0.3$ 88 $0.6$ $36.0$ $\mathbf{59.1\%}$ $\mathbf{78.8\%}$ $\mathbf{69.0\%}$ $\mathbf{52.7\%}$ $68.8\%$ $\mathbf{86.5\%}$
$0.6$ $0.3$ 88 $0.6$ $45.0$ $58.8\%$ $78.2\%$ $68.6\%$ $52.3\%$ $68.8\%$ $86.1\%$
$0.6$ $0.3$ 88 $0.6$ $60.0$ $58.4\%$ $77.9\%$ $67.5\%$ $51.3\%$ $\mathbf{69.6\%}$ $85.5\%$
$0.6$ $0.3$ 88 $0.6$ $90.0$ $56.2\%$ $76.6\%$ $64.8\%$ $48.5\%$ $68.3\%$ $84.7\%$
$0.6$ $0.3$ 88 $0.6$ $180.0$ $51.2\%$ $71.9\%$ $56.3\%$ $\% 64.264.2\%$ $82.9\%$
$0.6$ $0.3$ 66 $0.6$ $20.0$ $57.3\%$ $80.0\%$ $65.2\%$ $51.2\%$ $65.8\%$ $84.1\%$
$0.6$ $0.3$ 66 $0.6$ $27.5$ $59.4\%$ $81.3\%$ $68.5\%$ $53.4\%$ $68.8\%$ $\mathbf{85.7\%}$
$0.6$ $0.3$ 66 $0.6$ $30.0$ $59.7\%$ $\mathbf{81.7\%}$ $\mathbf{69.0\%}$ $\mathbf{53.7\%}$ $69.0\%$ $85.3\%$
$0.6$ $0.3$ 66 $0.6$ $36.0$ $\mathbf{59.9\%}$ $81.4\%$ $\mathbf{69.0\%}$ $53.6\%$ $\mathbf{69.8\%}$ $\mathbf{85.7\%}$
$0.6$ $0.3$ 66 $0.6$ $45.0$ $59.3\%$ $80.5\%$ $67.9\%$ $52.8\%$ $69.4\%$ $85.6\%$
$0.6$ $0.3$ 66 $0.6$ $60.0$ $58.5\%$ $79.6\%$ $67.0\%$ $51.5\%$ $69.6\%$ $84.9\%$
$0.6$ $0.3$ 66 $0.6$ $90.0$ $55.7\%$ $77.2\%$ $62.8\%$ $48.1\%$ $67.9\%$ $84.3\%$
$0.6$ $0.3$ 66 $0.6$ $180.0$ $48.8\%$ $70.6\%$ $50.9\%$ $40.8\%$ $61.1\%$ $83.4\%$
$0.6$ $0.6$ 66 $0.6$ $20.0$ $57.4\%$ $78.2\%$ $66.6\%$ $51.7\%$ $64.9\%$ $83.7\%$
$0.6$ $0.6$ 66 $0.6$ $27.5$ $60.0\%$ $79.4\%$ $70.1\%$ $53.5\%$ $68.4\%$ $\mathbf{86.1\%}$
$0.6$ $0.6$ 66 $0.6$ $30.0$ $60.0\%$ $\mathbf{79.7\%}$ $\mathbf{70.8\%}$ $\mathbf{53.6\%}$ $69.9\%$ $\mathbf{86.1\%}$
$0.6$ $0.6$ 66 $0.6$ $36.0$ $\mathbf{60.1\%}$ $79.6\%$ $\mathbf{70.8\%}$ $53.2\%$ $70.3\%$ $85.7\%$
$0.6$ $0.6$ 66 $0.6$ $45.0$ $59.5\%$ $79.0\%$ $69.5\%$ $52.5\%$ $70.8\%$ $85.3\%$
$0.6$ $0.6$ 66 $0.6$ $60.0$ $58.8\%$ $78.5\%$ $69.0\%$ $51.1\%$ $\mathbf{71.1\%}$ $85.3\%$
$0.6$ $0.6$ 66 $0.6$ $90.0$ $56.5\%$ $76.5\%$ $65.4\%$ $48.2\%$ $70.1\%$ $83.8\%$
$0.6$ $0.6$ 66 $0.6$ $180.0$ $50.9\%$ $71.4\%$ $56.2\%$ $41.6\%$ $65.4\%$ $82.9\%$
$0.8$ $0.8$ 66 $0.6$ $20.0$ $56.6\%$ $80.7\%$ $65.5\%$ $50.8\%$ $65.4\%$ $82.6\%$
$0.8$ $0.8$ 66 $0.6$ $27.5$ $58.7\%$ $81.9\%$ $68.9\%$ $\mathbf{52.8\%}$ $68.6\%$ $84.9\%$
$0.8$ $0.8$ 66 $0.6$ $30.0$ $\mathbf{59.1\%}$ $\mathbf{82.0\%}$ $\mathbf{69.2\%}$ $52.7\%$ $69.5\%$ $84.8\%$
$0.8$ $0.8$ 66 $0.6$ $36.0$ $58.9\%$ $81.7\%$ $68.8\%$ $52.6\%$ $\mathbf{69.9\%}$ $85.0\%$
$0.8$ $0.8$ 66 $0.6$ $45.0$ $58.4\%$ $81.4\%$ $67.7\%$ $51.6\%$ $69.8\%$ $\mathbf{85.1\%}$
$0.8$ $0.8$ 66 $0.6$ $60.0$ $57.4\%$ $80.1\%$ $66.6\%$ $50.2\%$ $\mathbf{69.9\%}$ $84.3\%$
$0.8$ $0.8$ 66 $0.6$ $90.0$ $55.7\%$ $78.6\%$ $63.7\%$ $47.8\%$ $68.9\%$ $84.2\%$
$0.8$ $0.8$ 66 $0.6$ $180.0$ $50.2\%$ $74.0\%$ $55.2\%$ $41.5\%$ $63.7\%$ $83.1\%$
A.4 Training & Inference Details
We train our models using mixed precision to optimize both memory efficiency and training speed. The training process spans 100,000100{,}000 iterations, utilizing the Adam optimizer [27] with a OneCycle learning rate schedule [44], which gradually decays from a peak learning rate to enhance convergence.
Consistent with RVT [18], we employ a mixed batching strategy to balance computational efficiency and memory usage. Specifically:
Our training process utilizes the YOLOX framework [63], a versatile object detection framework known for its efficient and high-performing architecture. We employ a multi-component loss function to optimize our model effectively:
To ensure stable training, these loss components are averaged across both the batch and sequence length at each optimization step. This averaging process helps to reduce variance during training and facilitates smoother convergence of the model parameters. We also include a regularization term to balance the contribution of both modalities.
Training Configuration. The training is conducted with a batch size of 8, which provides an optimal balance between efficient GPU utilization and memory requirements. Each training sample contains a sequence length of 11 frames, allowing the model to learn temporal dependencies effectively. The frame backbone’s weights are initialized using pre-trained ResNet, and the event backbone’s weights are initialized using pre-trained RVT. The learning rate is set to $1\times 10^{-4}$, following a OneCycle learning rate schedule that allows efficient exploration of the learning space and helps in achieving faster convergence.
Hardware & Training Time. All training experiments are carried out on two NVIDIA RTX A50005000 GPUs, each with 2424GB of memory, providing the computational resources necessary for handling the high-resolution event data and RGB frames. The complete training process, including all iterations and model optimization, takes approximately one day, demonstrating the efficiency of our implementation in terms of both training speed and resource utilization.
Appendix B Additional Quantitative Results
B.1 Complete Results of Efficiency Analysis
We include the complete results of the efficiency analysis in Tab. 8.
@@ B.2 Complete Results of Ablation Study
B.3 Complete Results of Hyperparameter Search
@@
Appendix C Additional Qualitative Results
C.1 Qualitative Comparisons of Event Detectors
We present additional qualitative comparisons in Fig. 9, Fig. 10, and Fig. 11 to further illustrate the advantages of FlexEvent over three state-of-the-art methods – RVT [18], SAST [39], and DAGr [16] – across nine diverse scenes.
As shown in Fig. 9, under fast-motion conditions, RVT and SAST fail to detect pedestrians, while DAGr misclassifies many distant objects as pedestrians or vehicles. In contrast, FlexEvent accurately captures all objects, demonstrating superior robustness in challenging high-speed scenarios.
Similarly, in Fig. 10, RVT frequently misses objects in cluttered scenes, and both SAST and DAGr mistakenly recognize pedestrians and distant vehicles under noisy, occluded conditions. Here again, FlexEvent excels by reliably detecting all relevant objects, highlighting its ability to handle complex urban environments.
Finally, in Fig. 11, rapidly changing illumination leads RVT and DAGr to misinterpret subtle motion cues; yet FlexEvent preserves stable performance, underscoring its enhanced temporal awareness. These observations confirm that FlexEvent not only surpasses competing methods in quantitative benchmarks but also consistently delivers high detection accuracy in real-world conditions – ensuring reliable performance in fast-moving, cluttered, or dynamically lit scenes.
C.2 Comparisons under Different Frequencies
We further evaluate the performance of FlexEvent across diverse event frequencies in Fig. 12, Fig. 13, and Fig. 14. Specifically, we compare our method with RVT under settings ranging from low-frequency ($20$ Hz) to high-frequency ($180$ Hz) event streams, including an extreme scenario with empty events.
In Fig. 12, RVT struggles to detect distant or partially occluded objects at both low and high frequencies, whereas FlexEvent consistently identifies all targets by adaptively fusing frame-based evidence with event cues.
Similarly, in Fig. 13, RVT suffers substantial performance drops under sparse event conditions, often overlooking pedestrians; yet FlexEvent maintains reliable detection through its frequency-adaptive training strategy.
Finally, in Fig. 14, even when event input is minimal or missing, FlexEvent retains high accuracy by leveraging complementary frame information. These comparisons reaffirm our robustness and adaptability, highlighting our ability to handle a broad spectrum of event frequencies while preserving superior detection performance.
Appendix D Potential Societal Impact & Limitations